A company manufactures two types of DVD players, A and B. Although there are differences among the products, there are a number of common parts within each player. Each model A requires one subassembly C and two subassembly D’s. Each subassembly C requires two F’s and two G’s. Each subassembly D requires two units of F. Each F requires two units of raw material I, while each G requires one unit of raw material H. Each model B requires two subassemblies C’s and two subassemblies E’s. Each E requires two units of G. The lead times and current inventory on hand of the parts and components are as follows:

Item

Number currently in stock

Leadtime (weeks)

DvD Model A

30

1

DvD Model B

50

2

Subassembly C

75

1

Subassembly D

80

2

Subassembly E

100

1

Part F

150

1

Part G

40

1

Raw material H

200

2

Raw material I

300

2

.

The Company has created a forecast that it plans to use as its master production schedule. Part of the MPS shows demand for 700 units of model A and 1200 units of model B in week ten. (a) Draw product trees for A and B. (b) Develop an MRP schedule for the products and each component to meet demand for DVD models A and B in week ten. Use chase strategy lot sizing.
Answer #1

(a)

1- 0A 1.3 (12) HI) 13

| L. 画画,

(b)

Use the following method to complete the MRP tables:

  1. Gross requirements of period-j (Gj) = Σk nk* (POReleasej of the parent-k) where nk = number of units of the item used in 1 unit of its parent-k
  2. Scheduled receipts in period-j (SRj): to be entered manually
  3. Projected on-hand in period-j (Ij) = Max.(0, Ij-1 + SRj + POReceiptj - Gj)
  4. Net requirements in period-j (Nj) = Max.(0, Gj - SRj - Ij-1)
  5. Planned order receipts of period-j (POReceiptj): to be entered manually when there is a positive net requirement based on the lot size rule.
  6. Planned order release of period-j (POReleasej) = POReceiptj+L
Item - A 1 2 3 4 5 6 7 8 9 10
Gross requirements 0 0 0 0 0 0 0 0 0 700
Scheduled receipts 0 0 0 0 0 0 0 0 0 0
Projected on-hand 30 30 30 30 30 30 30 30 30 30 0
Net requirements 0 0 0 0 0 0 0 0 0 670
Planned order release 0 0 0 0 0 0 0 0 0 670
Planned order receipt 0 0 0 0 0 0 0 0 670 0
Item - B 1 2 3 4 5 6 7 8 9 10
Gross requirements 0 0 0 0 0 0 0 0 0 1200
Scheduled receipts 0 0 0 0 0 0 0 0 0 0
Projected on-hand 50 50 50 50 50 50 50 50 50 50 0
Net requirements 0 0 0 0 0 0 0 0 0 1150
Planned order release 0 0 0 0 0 0 0 0 0 1150
Planned order receipt 0 0 0 0 0 0 0 1150 0 0
Item - C 1 2 3 4 5 6 7 8 9 10
Gross requirements 0 0 0 0 0 0 0 2300 670 0
Scheduled receipts 0 0 0 0 0 0 0 0 0 0
Projected on-hand 75 75 75 75 75 75 75 75 0 0 0
Net requirements 0 0 0 0 0 0 0 2225 670 0
Planned order release 0 0 0 0 0 0 0 2225 670 0
Planned order receipt 0 0 0 0 0 0 2225 670 0 0
Item - D 1 2 3 4 5 6 7 8 9 10
Gross requirements 0 0 0 0 0 0 0 0 1340 0
Scheduled receipts 0 0 0 0 0 0 0 0 0 0
Projected on-hand 80 80 80 80 80 80 80 80 80 0 0
Net requirements 0 0 0 0 0 0 0 0 1260 0
Planned order release 0 0 0 0 0 0 0 0 1260 0
Planned order receipt 0 0 0 0 0 0 1260 0 0 0
Item - E 1 2 3 4 5 6 7 8 9 10
Gross requirements 0 0 0 0 0 0 0 2300 0 0
Scheduled receipts 0 0 0 0 0 0 0 0 0 0
Projected on-hand 100 100 100 100 100 100 100 100 0 0 0
Net requirements 0 0 0 0 0 0 0 2200 0 0
Planned order release 0 0 0 0 0 0 0 2200 0 0
Planned order receipt 0 0 0 0 0 0 2200 0 0 0
Item - F 1 2 3 4 5 6 7 8 9 10
Gross requirements 0 0 0 0 0 0 6970 1340 0 0
Scheduled receipts 0 0 0 0 0 0 0 0 0 0
Projected on-hand 150 150 150 150 150 150 150 0 0 0 0
Net requirements 0 0 0 0 0 0 6820 1340 0 0
Planned order release 0 0 0 0 0 0 6820 1340 0 0
Planned order receipt 0 0 0 0 0 6820 1340 0 0 0
Item - G 1 2 3 4 5 6 7 8 9 10
Gross requirements 0 0 0 0 0 0 8850 1340 0 0
Scheduled receipts 0 0 0 0 0 0 0 0 0 0
Projected on-hand 40 40 40 40 40 40 40 0 0 0 0
Net requirements 0 0 0 0 0 0 8810 1340 0 0
Planned order release 0 0 0 0 0 0 8810 1340 0 0
Planned order receipt 0 0 0 0 0 8810 1340 0 0 0
Item - H 1 2 3 4 5 6 7 8 9 10
Gross requirements 0 0 0 0 0 8810 1340 0 0 0
Scheduled receipts 0 0 0 0 0 0 0 0 0 0
Projected on-hand 200 200 200 200 200 200 0 0 0 0 0
Net requirements 0 0 0 0 0 8610 1340 0 0 0
Planned order release 0 0 0 0 0 8610 1340 0 0 0
Planned order receipt 0 0 0 8610 1340 0 0 0 0 0
Item - I 1 2 3 4 5 6 7 8 9 10
Gross requirements 0 0 0 0 0 13640 2680 0 0 0
Scheduled receipts 0 0 0 0 0 0 0 0 0 0
Projected on-hand 300 300 300 300 300 300 0 0 0 0 0
Net requirements 0 0 0 0 0 13340 2680 0 0 0
Planned order release 0 0 0 0 0 13340 2680 0 0 0
Planned order receipt 0 0 0 13340 2680 0 0 0 0 0

Calculations in Excel (week 3 to 9 are hidden and contains formulae to be simply copied)

M 10 0 0 700 Item - A Gross requirements Scheduled receipts Projected on-hand Net requirements Planned order release Planned10 =2*M8 2*D8 =2*E8 Item -D Gross requirements Scheduled receipts Projected on-hand Net requirements Planned order release Pl

=2*(D40+D24) =2*(E40+E24) 10 =2* (M40+M24) Item -G Gross requirements Scheduled receipts Projected on-hand Net requirements P